# When Worse is Better: Navigating the compression-generation tradeoff in visual tokenization

### v1

@@ Abstract
Current image generation methods, such as latent diffusion and discrete token-based generation, depend on a two-stage training approach. In stage 1, an auto-encoder is trained to compress an image into a latent space; in stage 2, a generative model is trained to learn a distribution over that latent space. Most work focuses on maximizing stage 1 performance independent of stage 2, assuming better reconstruction always leads to better generation. However, we show this is not strictly true. Smaller stage 2 models can benefit from more compressed stage 1 latents even if reconstruction performance worsens, showing a fundamental trade-off between compression and generation modeling capacity. To better optimize this trade-off, we introduce Causally Regularized Tokenization (CRT), which uses knowledge of the stage 2 generation modeling procedure to embed useful inductive biases in stage 1 latents. This regularization makes stage 1 reconstruction performance worse, but makes stage 2 generation performance better by making the tokens easier to model: we are able to improve compute efficiency 2-3$\times$ over baseline and match state-of-the-art discrete autoregressive ImageNet generation (2.18 FID) with half the tokens per image (256 vs. 576) and a fourth the total model parameters (775M vs. 3.1B) as the previous SOTA.
@@ 1 Introduction
Modern image generation methods use a two-stage approach to training. In the first stage, a model such as a VQGAN [9] (for autoregressive modeling e.g., [54]) or a continuous VAE (for diffusion [40]) is trained to compress images to a latent representation. In the second stage, a generative model is trained to learn a prior over the latent distribution induced by the first stage. Generation is done by using the second stage model to sample a latent, and then applying the decoder of the first stage model to map back to image space.
@@
Figure 1: Overview of our method. We introduce a small 2-layer causal transformer [35, 46, 8] trained to optimize $\ell_{2}$ next-token prediction on the pre-quantized latents of the auto-encoder. This loss is propagated through the encoder, producing tokens with a causal transformer inductive bias. Thus, we call our method Causally Regularized Tokenization (CRT).
This setup reveals an interesting interaction between the stage 1 and stage 2 models. The most important interacting variables for generation performance are: (1) The reconstruction performance of the stage 1 auto-encoder (its “distortion”), (2) The degree of compression of the stage 1 auto-encoder (its “rate”) [^1] Note that the concept of “rate” is the entropy of the distribution, so the theoretical lower bound on compression, not the compression ratio we actually achieve. and (3) The smallest achievable loss of the stage 2 model in learning a distribution over this compressed representation (the “irreducible loss” of distribution modeling). Much empirical work has gone into improving (1), as it is generally seen as the lower bound of generation performance. Indeed, scaling laws [15, 23, 59, 19] indicate that stage 2 models can be made arbitrarily powerful and are therefore bottlenecked by stage 1 model performance. But what if our stage 2 model is limited (e.g., by inference time compute limitations)? Given a limited capacity stage 2 model, what is the optimal stage 1 auto-encoder? We view this as a trade-off between rate, distortion, and distribution modeling. Reducing the rate also increases the distortion but would improve distribution modeling performance because the irreducible loss is lower-bounded by rate. We hypothesize that optimizing this trade-off between distortion and distribution modeling would yield substantial generation improvements and enable parameter-efficient stage 2 training.
@@
We study this in the context of a vector-quantized GAN (VQGAN) stage 1 model and an auto-regressive stage 2 model trained on VQ tokens. We explore key interventions which affect the stage 1 rate: (a) Number of tokens per image, (b) Dictionary size of the VQ model, and (c) Data scaling, and examine how they affect stage 2 generation performance as a function of model scale. We show that small models benefit strongly from more compressed representations, and that this relationship also depends on stage 2 data scale. Given these findings, we ask: Can we design an optimal stage 1 tokenizer given that we know our stage 2 model is auto-regressive? Through a simple model-based causal loss, we learn a causal inductive bias on our stage 1 tokenizer. Intuitively, our loss attempts to make token $i$ as predictable as possible given tokens 0 through $i-1$. We call this training recipe Causally Regularized Tokenization (CRT). When equal training compute is used for stage 1, this regularization makes reconstruction performance worse, but improves stage 2 generation performance across ImageNet [41] and LSUN [52]. With CRT, we also show significant improvements in stage 2 model performance scaling, especially for small models.
Summary of major contributions.
@@
We study the complex trade-off between compression and generation by varying codebook size, tokens per image, and stage 2 training compute. Our analysis shows that the ideal amount of image compression varies with generation model capacity, challenging conventional wisdom about reconstruction quality.
@@
$\cdot$ We introduce a method for training a stage 1 tokenizer with a causal inductive bias. This improves inference and training compute scaling of our stage 2 models, without any other interventions, leading to our key result: by making tokens easier to model, we improve compute efficiency 2-3x over baseline. We also match SOTA discrete autoregressive ImageNet generation (2.18 FID) with half the tokens per image (256 vs 576) and a third the total model parameters (740M vs. 3.1B) as the previous SOTA model [43].
@@ 2 Related Work
Evolution of visual tokenization. The current paradigm of discrete visual tokenizers came from VQ-VAE [49], originally introduced as a method for discrete latent representation learning in continuous domains such as image and audio. It was introduced to prevent the posterior collapse problem in VAEs by separating the compression and generation phases of training. Section 4.2 of [49] notes that their loss function leads to blurry reconstructions, and the authors suggest including perceptual metrics (metrics that align with human judgment of similarity) in the loss function. Esser et al. [10] explores this idea with VQGAN, which introduces a perceptual loss based on $LPIPS$ [58], and a GAN loss to add realistic textures. Since the advent of VQGAN, there have been numerous attempts at improving visual tokenizers [53, 14, 61, 60, 57, 50, 22, 39, 51, 34, 62, 38]. Some works focus on improvements including ViT-VQGAN [53], which replaces CNN layers with ViT layers, and MoVQGAN [61], which adds adaptive instance normalization layers in the CNN decoder. Other works focus on mitigating “codebook collapse” [20] in vector quantization (VQ) [13] by introducing novel quantization schemes, for example by replacing dead codes with encoder outputs [7]. We refer the reader to [20] for a thorough review of these methods. Other works replace VQ entirely with scalar quantization techniques [32], and demonstrates increased codebook utilization compared to VQ: for example, MAG-VIT [55, 31] quantizes each dimension to 2 values and adds entropy loss to encourage codebook utilization.
Trade-offs between stage 1 and stage 2 performance. We are not the first to notice a disconnect between stage $1$ and stage $2$ performance. [47] extensively studies how different regularization methods affect the usefulness of learned VAE representations for downstream representations, introducing the “rate-distortion-usefulness” trade-off. Notably, they highlight that stage 1 loss (e.g. VAE loss) is not representative of downstream task usefulness. [48] also shows in Figure 4 demonstrates that increasing $\beta$ in their $\beta$-VAE [17] architecture worsens reconstruction but improves generation performance. Recent advances in lookup-free quantization [32, 55] have unlocked better codebook utilization, improving stage 1 performance. Both works observe that scaling up codebook size improves stage 1 performance but degrades stage 2 performance (see Figure 3 in [32] and Figure 1 in [55]). Outside of visual tokenization, previous work in text tokenization observes that optimizing text tokenizer compression (i.e. stage 1 performance) leads to worse perplexity [28] and downstream accuracy [42].
@@ 3 Experiments & Methodology
Structure of our study. Our goal is to understand the variables involved in tokenizer construction affecting generation performance at different model scales. First, fixing a stage 1 tokenizer, we rigorously study the connection between generation performance and compute scaling, establishing gFID on ImageNet as a consistent generation metric to center our study. [^2] This consistency is not as present when using gFID on non-ImageNet datasets because the Inception network [44] is in-distribution on ImageNet. At the core of this exploration is the rate-distortion trade-off in the stage 1 tokenizer. How does changing the stage 1 “rate” (bits per output pixel) affect stage 2 model performance at various scales? We study scaling properties with respect to key tokenizer interventions which affect this rate-distortion trade-off. Finally, we study a new angle on the rate-distortion which affects compute-optimal generation performance scaling: causally regularized tokenization (CRT).
@@ 3.1 VQGAN and Auto-regressive Image Generation
We summarize the common auto-regressive image generation procedure we use for the basis of our study [54, 43, 9].
Stage 1 (Tokenization). In this stage, the goal is to map an image $x\in\mathbb{R}^{H\times W\times 3}$ to a set of $N$ discrete tokens $X_{i}\in\mathcal{C}$, where $\mathcal{C}$ is a codebook such that $\mathcal{C}\subset\mathbb{R}^{d}$. We also learn a decoder which can reconstruct the original image from a set of tokens. We use the VQGAN architecture [9], which is a ResNet-based architecture that is a mix of convolutions and self-attention layers in both the encoder and decoder. During training, the encoder maps the input image $x$ to continuous latents $\hat{X}_{i}$. Then, we use a codebook look-up to determine the nearest embedding $X_{i}\in\mathcal{C}$. Finally, these are passed to the decoder to get a reconstruction.
Figure 3: Neural scaling laws for auto-regressive image generation (ImageNet 256x256). We show validation cross-entropy loss (left) and gFID (center) as a function of total train FLOPs. Across model scales, we observe log-log linear scaling with train FLOPs for both FID and cross-entropy loss across Pareto frontiers. Moreover, we observe validation cross entropy loss is generally predictive of gFID (right). However, at large model scales (yellow-line), we notice that over-training can lead increase in validation cross-entropy but decrease in gFID, implying that validation loss is predictive of gFID only when training duration is held consistent (see Section 3.3.1).
Stage 2 (Generation). In this stage, the goal is to learn a prior over the latent tokens learned in the previous stage. For each image in our train set, we use the encoder learned in stage 1 to map it to a set of discrete tokens. We fix an order with which to decode these tokens per image (raster-scan works best empirically), and treat these as sequences. Finally, we train a transformer with causal attention to learn the conditional distributions $p(x_{i}|x_{0}\ldots x_{i-1})$ for all tokens $x_{i}$ by maximizing $\prod_{i=0}^{N}q_{\theta}(x_{i}|x_{0}\ldots x_{i-1})$ for our model $q_{\theta}$. This is done by minimizing cross-entropy loss over the next-token prediction objective. We train our generation models in the class-conditional regime. To do this, we append a learned class token to the start of the sequence. We also use classifier-free guidance (CFG) at inference time, where we predict a token based on both the class-conditional and unconditional distributions. These decisions are discussed in detail in Section 3.6. Details on architecture, training, and CFG discussed in Appendix A.
@@ 3.2 Metrics
The two categories for evaluation are reconstruction performance (to evaluate the stage 1 model) and generation performance (to evaluate the stage 2 + stage 1 pipeline).
Reconstruction metrics. We use Fréchet Inception Distance (FID) [16] between validation and a reconstructed validation set as our primary metric. We refer to this metric as rFID (reconstruction FID). We also report the signal-to-noise ratio (PSNR) and the multi-scale structured similarity metric (MS-SSIM) [1].
Generation metrics. We use FID between the generations and the whole validation set for our primary metric. For clarity, we refer to this as generation FID (gFID). To conform with existing literature, we use the OpenAI guided diffusion [6] repository to evaluate our models. We report Inception Score, precision, and recall, which are proxies for generation fidelity and diversity.
@@ 3.3.1 Neural Scaling Laws for Image Generation
Setup. We fix a stage 1 tokenizer (16k token codebook size, 256 sequence length), vary stage 2 model parameters and compute dedicated to training. Fixing codebook size is important because, all else being equal, validation cross-entropy loss is proportional to the intrinsic entropy of the codes, which increases log-log-linearly by codebook size (see Figure 10). We only plot the end point of each training run since our cosine learning rate decay schedule results in substantial validation loss changes in the last few iterations of training, yielding inconsistent scaling trends. Given that we are iterating on a fixed-size dataset (about $3\times 10^{8}$ total tokens at 256 tokens per image), we do not scale models past 740M parameters to avoid data scale bottlenecks.
@@
Log-log-linear scaling for image generation. In Figure 3, we present our findings on the relationship between compute scaling and generation performance. These results corroborate scaling trends found in [15, 23]. We see consistent log-log-linear scaling with validation loss (Figure 3 (left)) on the compute validation loss Pareto curve across four orders of magnitude. These validation scaling loss trends serve as the basis for our scaling studies.
@@
Validation loss and generation performance. In Figure 3 (center), we plot the gFID score as a function of training compute for different model scales. We see that, in general, lower validation loss implies better generation performance (lower gFID). However, if we look at the upper Pareto frontier for Figure 3 (left) (highest gFID score for lowest validation loss), we see that individual model scales achieve lower gFID for correspondingly higher validation losses. This is considered the “overtraining” regime with respect to validation loss, described in [12]. This implies that validation loss is only predictive of FID when the number of training steps is held constant. Therefore, compute optimality with respect to validation loss also corresponds to compute optimality with respect to gFID, meaning overall scaling bounds transfer, even if specific trends do not.
Figure 4: Relationship between tokens per image and model size. We show compute scaling laws for stage 2 performance for different stage 1 tokens/image. Each plot uses a different stage 2 model size (left: 110M, center: 340M, right: 550M) trained across several durations (separate training runs). We notice that larger models can take advantage of more tokens/image, whereas smaller models are more compute efficient fewer tokens.
Figure 5: Performance improves reliably with increase tokens/image. We demonstrate the results of a simple sweep of tokens/image across our three main reconstruction metrics. Our base model is trained with 256 tokens/image (16k codebook size), and we increase the tokens/image at inference time using the method described in Section 3.4. We observe that increasing tokens/image consistently improves reconstruction performance. We also show the performance of CRT tokens, which has the worse (reconstruction) but better (generation) phenomena.

### v2

# When Worse is Better: Navigating the Compression-Generation Trade-off for Visual Tokenization

@@ Abstract
[…]
@@ 1 Introduction
Modern image generation methods use a two-stage approach to training. In the first stage, a model such as a VQGAN [esser2020taming, yu2022scaling, rombach2021high] is trained to compress images to a latent representation. In the second stage, a generative model is trained on these latent representations. This setup implies a deep interaction between the stage 1 and stage 2 models. Since the stage 1 model does not have to reconstruct the image perfectly, it can compress the image more aggressively to achieve a simpler latent distribution for the stage 2 model to learn. However, we ultimately care about the quality of the generated images, which depends on both the stage 1 model’s ability to reconstruct and the stage 2 model’s ability to learn the latent distribution. On one extreme, if the stage 1 model compresses to a constant, the resulting latent distribution is easy to model but can only generate one image. At the other extreme, the stage 1 model does not compress at all and we get no benefit from latent distribution modeling.
We view this as a trade-off between rate, distortion, and distribution modeling, similar to the rate-distortion-usefulness trade-off discussed in [tschannen2018recent]. The important interacting variables for generation performance are: (1) The reconstruction performance of the stage 1 auto-encoder (its “distortion”), (2) The degree of compression of the stage 1 auto-encoder (its “rate”) and (3) The smallest achievable loss of the stage 2 model in learning a distribution over this compressed representation (the “irreducible loss” of distribution modeling).
@@
Figure 1: Overview of our method. We introduce a small 2-layer causal transformer [radford2018improving, touvron2023llama, el2024scalable] trained to optimize $\ell_{2}$ next-token prediction on the pre-quantized latents of the auto-encoder. This loss is propagated through the encoder, producing tokens with a causal transformer inductive bias. Thus, we call our method Causally Regularized Tokenization (CRT).
Figure 2: Qualitative examples from our best model (2.18 gFID). We show our model’s ability to generate high-quality, diverse images on par with LlamaGen-3B [sun2024autoregressive] with only 775M parameters and <<50% of the tokens per image. For more non-curated examples see Figure 14 and Appendix G.
@@
We study these factors using a vector-quantized GAN (VQGAN) stage 1 model and an auto-regressive stage 2 model trained on VQ tokens. We explore key interventions which affect the stage 1 rate (bits per image pixel) and examine how they affect stage 2 generation performance as a function of model scale. Through the lens of compute scaling laws, we show that compression is not all you need — while small models benefit from more compressed representations, and this relationship also depends on stage 2 training compute. Given this, we ask: can we design an optimal stage 1 tokenizer given that our stage 2 model is auto-regressive? Through a simple model-based causal loss, we learn a causal inductive bias on our stage 1 tokenizer. Intuitively, our loss attempts to make token $i$ as predictable as possible given tokens 0 through i1i-1. We call this training recipe Causally Regularized Tokenization (CRT). This regularization trades off reconstruction for distribution modeling, improving generation performance across ImageNet [russakovsky2014imagenet] and LSUN [yu2015lsun]. With CRT, we also show significant improvements in generation performance scaling laws (2-2.5x faster training). We demonstrate our improvements through comprehensive experiments across five orders of magnitude of training compute and two orders of magnitude of model scale.
Summary of contributions.
@@
We study the complex trade-off between compression and generation. Our analysis shows that the ideal amount of image compression varies with generation model capacity.
@@
$\cdot$ We introduce a method for training a stage 1 tokenizer with a causal inductive bias. This improves inference and training compute scaling of our stage 2 models, without any other interventions, leading to our key result: by making tokens easier to model, we improve compute efficiency 2-3x over baseline. We also match LlamaGen-3B [sun2024autoregressive], a prior SOTA auto-regressive discrete generation model (2.18 FID) with half the tokens per image (256 vs 576) and a fourth the total model parameters (775M vs. 3.1B).
@@ 2 Related Work
For an extensive rundown of visual tokenization methods, we refer the reader to Section I.
Trade-offs between stage 1 and stage 2 performance.: We are not the first to notice a disconnect between stage $1$ and stage $2$ performance. [tschannen2018recent] extensively studies how different regularization methods affect the usefulness of learned VAE representations for downstream representations, introducing the “rate-distortion-usefulness” trade-off. Notably, they highlight that stage 1 loss (e.g. VAE loss) is not representative of downstream task usefulness. [tschannen2025givt] also shows in Figure 4 demonstrates that increasing $\beta$ in their $\beta$-VAE [higgins2017beta] architecture worsens reconstruction but improves generation performance. Recent advances in lookup-free quantization [mentzer2023finite, yu2023language] achieve better codebook utilization, improving stage 1 performance. Both works observe that scaling up codebook size improves stage 1 performance but degrades stage 2 performance (see Figure 3 in [mentzer2023finite] and Figure 1 in [yu2023language]). Outside visual tokenization, previous work in text tokenization observes that optimizing text tokenizer compression (i.e. stage 1 performance) leads to worse perplexity [lester2024training] and downstream accuracy [schmidt2024tokenization]. JetFormer [tschannen2024jetformer] is an end-to-end training procedure for generation, using an auto-regressive (AR) model in conjunction with a flow-based model to perform generation and compression at the same time. They balance this trade-off by using the flow-based model for compression, which by its invertible nature, does not allow for information collapse.
1D-tokenizers for Generative Visual Models. Many recent works have identified the issue that existing visual tokenizers are not built for auto-regressive generation and seek to rectify this issue through the use of so-called “1D” tokenizers, which focus on global over local semantics. [ge2023plantingseedvisionlarge] constructs a discrete tokenizer that captures high-level semantics by using Stable Diffusion [rombach2021high] as a stage 2 model. VAR and VQVAE–2 [tian2024visual, razavi2019generating] modifies the tokenizer to produce multi-scale tokens for easier auto-regressive modeling. [yu2024image] and [wang2024larptokenizingvideoslearned] use a transformer register-based [darcet2023vision] approach to learn global tokens with perceptual losses. SEED [ge2307planting] constructs a 1D tokenizer by using Stable Diffusion [rombach2021high] as a powerful decoder on quantized tokens. More recently, methods [bachmann2025flextok, wen2025principal, miwa2025one] extend this strategy to create progressive tokens, which can be used to flexibly refine generated images for sampling efficiency. In contrast to these works, we do not modify the common VQGAN tokenizer architecture and focus on constructing a causal regularization that enforces compatibility with a stage 2 auto-regressive model. This makes our work generally applicable, as variants of this architecture are common across modern generation systems. We further rigorously study scaling properties with respect to this regularization and other common interventions on tokenizer compression rate, which is not found in existing literature.
@@ 3 Experiments & Methodology
Structure of our study. Our goal is to understand the variables involved in tokenizer construction affecting generation performance at different model scales. First, we fix a stage 1 tokenizer, and study the connection between generation performance (gFID on ImageNet) and compute scaling. We study how changing the stage 1 “rate” (bits per output pixel) affect stages 2 model performance at various scales. Then we examine a new angle on the rate-distortion which affects compute-optimal generation performance scaling: causally regularized tokenization (CRT).
@@ 3.1 VQGAN and Auto-regressive Image Generation
We summarize the standard auto-regressive image generation procedure [yu2022scaling, sun2024autoregressive, esser2020taming].
Stage 1 (Tokenization). In this stage, the goal is to map an image $x\in\mathbb{R}^{H\times W\times 3}$ to a set of $N$ discrete tokens $X_{i}\in\mathcal{C}$, where $\mathcal{C}$ is a codebook such that $\mathcal{C}\subset\mathbb{R}^{d}$. We also learn a decoder which can reconstruct the original image from a set of tokens. We use the VQGAN architecture [esser2020taming], which is a ResNet-based architecture that is a mix of convolutions and self-attention layers in both the encoder and decoder. During training, the encoder maps the input image $x$ to continuous latents $\hat{X}_{i}$. Then, we use a codebook look-up to determine the nearest embedding $X_{i}\in\mathcal{C}$. Finally, these are passed to the decoder to get a reconstruction.
Stage 2 (Generation). In this stage, the goal is to learn a prior over the latent tokens learned in the previous stage. For each image in our train set, we use the encoder learned in stage 1 to map it to a set of discrete tokens. We fix an order with which to decode these tokens per image (raster-scan works best empirically [esser2020taming]), and treat these as sequences. Finally, we train a transformer with causal attention to learn the conditional distributions $p(x_{i}|x_{0}\ldots x_{i-1})$ for all tokens $x_{i}$ by maximizing $\prod_{i=0}^{N}q_{\theta}(x_{i}|x_{0}\ldots x_{i-1})$ for our model $q_{\theta}$. This is done by minimizing cross-entropy loss over the next-token prediction objective. We train our generation models in the class-conditional regime. To do this, we append a learned class token to the start of the sequence. We also use classifier-free guidance (CFG) at inference time, where we predict a token based on both the class-conditional and unconditional distributions. These decisions are discussed in detail in Section 3.6. Details on architecture, training, and CFG discussed in Appendix A.
@@ 3.2 Metrics
Reconstruction metrics. We use Fréchet Inception Distance (FID) [heusel2017gans] on a validation set as our primary metric. We refer to this metric as rFID (reconstruction FID). We also report the signal-to-noise ratio (PSNR) and multi-scale structured similarity metric (MS-SSIM) [wang2003multiscale].
Generation metrics. We use FID between the generations and the whole validation set for our primary metric. For clarity, we refer to this as generation FID (gFID). To conform with existing literature, we use the OpenAI guided diffusion [dhariwal2021diffusion] repository to evaluate our models. We also report Inception Score, precision, and recall, which are proxies for generation fidelity and diversity.
@@ 3.3.1 Neural Scaling Laws for Image Generation
Scaling Laws for Generation Performance. We want to study the relationship between validation loss scaling laws (as studied in [kaplan2020scaling] and [henighan2020scaling]) and generation performance scaling as measured by FID. For modeling scaling laws, we use the common [kaplan2020scaling, hoffmann2022training] functional form $L(C)=L_{min}+L^{\prime}(C)$ where $L_{min}$ is considered the irreducible loss of the problem and $L^{\prime}$ is the reducible loss as a function of $C$ (compute). In our case, we fit $L^{\prime}(C)=C^{\alpha}e^{\lambda}$. When graphed on a log-log scale, $\alpha$ is the slope and $\lambda$ is the $y$-intercept of a line. For both gFID and validation loss scaling laws, we fit $L_{min}$ empirically. We note that rFID is usually equivalent to $L_{min}$, however this is not always the case in practice (see Figure 6 (far left)).
@@
Experimental Setup. We fix a stage 1 tokenizer (16k token codebook size, 256 sequence length), vary stage 2 model parameters and compute dedicated to training. Fixing codebook size is important because, all else being equal, validation cross-entropy loss is proportional to the intrinsic entropy of the codes, which increases log-log-linearly by codebook size (see Figure 9). Each point in these scaling law plots is an independent training run. We only plot the end point of each training run since our cosine learning rate decay schedule results in substantial validation loss changes in the last few iterations of training. Given that we are iterating on a fixed-size dataset ($\sim 3\times 10^{8}$ total tokens), we do not scale models past 775M parameters to avoid data scale bottlenecks.
@@
Log-log-linear loss scaling for visual token modeling. In Figure 4, we present our findings on the relationship between compute scaling and validation loss. These results corroborate scaling trends found in [henighan2020scaling, kaplan2020scaling]. We see consistent log-log-linear scaling with validation loss (Figure 4 (left)) on the compute validation loss Pareto curve across four orders of magnitude.
Figure 3: Relationship between different methods of scaling bits/pixel and reconstruction performance. Here we increase bits per pixel of either vary the codebook size $|C|\in\{2^{10}\ldots 2^{17}\}$ with fixed number of tokens $N=256$ or by varying $N\in\{196\ldots 576\}$ with fixed $|C|=2^{14}$ using the strategy described in Section 3.4. Increasing token count is generally more effective at the cost of increased inference compute. We also compare our VQ recipe with FSQ [mentzer2023finite], and find it generally outperforms, so we use VQ for all of our experiments going forward.
Figure 4: Neural scaling laws for auto-regressive image generation (ImageNet 256x256). We show validation cross-entropy loss (left) and gFID (center) as a function of total train FLOPs. Across model scales, we observe log-log linear scaling with train FLOPs for both FID and cross-entropy loss across Pareto frontiers. Moreover, we observe validation cross entropy loss is generally predictive of gFID (right). See Section 3.3.1 for further discussion.
Figure 5: Relationship between compute scaling laws and tokens per image. We show compute scaling laws for stage 2 performance for different stage 1 tokens/image. We use the same base tokenizer, expanded to greater tokens/image counts using the strategy described in Section 3.4. We see that fewer tokens per image is generally more compute efficient until we get close to saturation (rFID lower bound), despite their much better reconstruction performance. Also, (far right) shows that the true trend deviates more from the predicted trend as tokens per image count increases close to saturation, implying larger models are needed to maintain compute optimality.
Figure 6: Relationship between compute scaling laws and stage 1 codebook size. We show individual model scaling laws at three different stage 1 codebook sizes. 1k outperforms 131k at low compute regimes and the trend reverses at high compute regimes. Note that 16k outperforms both even though it has worse rFID that the 131k codebook size tokenizer (2.21 rFID vs 1.72 rFID).
@@
Validation loss and generation performance. In Figure 4 (center), we plot the gFID score as a function of training compute for different model scales. We see that, in general, lower validation loss implies better generation performance (lower gFID). However, if we look at the upper Pareto frontier for Figure 4 (right), we see that individual model scales can achieve lower gFID for correspondingly higher validation losses. This is considered the “overtraining” regime with respect to validation loss, described in [gadre2024language]. This implies that validation loss is only predictive of FID ordering when the number of training steps is held constant. More generally, models on the training compute pareto frontier of validation loss are also on the pareto frontier of gFID. This means that compute optimality with respect to validation loss also corresponds to compute optimality with respect to gFID, even if specific orderings are not respected between the two quantities.